OpenStudy (anonymous):

Summation question: How do I find the sum of [(i-1)^2 + (i+3)^3] when i= 1 and n= 4? Thanks in advance for help!

OpenStudy (anonymous):

\[\sum_{i=1}^{4} [(i-1)^2 +(i+3)^3]\]

OpenStudy (anonymous):

is i real or imaginary?

OpenStudy (anonymous):

i is the index of summation

OpenStudy (anonymous):

I is the variable that is changing in the summation

OpenStudy (anonymous):

so you can do this step by step, start with i= 1, (1-1)^2+(1+3)^2 then you can move to i=2, (2-1)^2+(2+3)^2 and then repeat till i = 4, and add all of them together.

OpenStudy (anonymous):

by i =3 you should see a pattern.

OpenStudy (anonymous):

ok so i just find the answer for each i, of which there are 4, and add them up? my brain makes out summations to be harder than they are, i suppose. but for your second part, the second parentheses is cubed

OpenStudy (anonymous):

(i+3)^3